OpenStudy (anonymous):

please help me!!! Solve. 5 = –(–z + 3) A.–8 B.8 C.–2 D.2

OpenStudy (paki):

5 = –(–z + 3) 5= -z+3 z=3-5 z=-2

OpenStudy (anonymous):

thank you!!!

OpenStudy (paki):

pleasure

OpenStudy (anonymous):

can you help with one more?

OpenStudy (paki):

yeah